Greenhab Report – April 15th

GreenHab Officer: David Murray

Environmental control: Ambient with door open 9:00 to ? (high winds blew the door shut at some point while we were gone)

Shade Cloth: 80%

Low Temperature: 15 C

High Temperature: 37 C

Relative humidity: 17%

Hours of Supplemental Light: 0

Daily Water Usage for Crops: 11 L

Water in Blue Tank: Low, needs refilling

Time(s) of Watering: 11:40

Changes to Crops: None.

Narrative: Crops are looking good but the tomatoes are starting to show signs of nutrient deficiency. I’ll fertilize them with Miracle Grow tomorrow and see if it improves. It’s been hot, but the fan is running on it’s highest setting throughout the hottest parts of the day. The GreenHab door shut at some point before 6:45 due to high winds.

Harvest: None

Operations Report – April 15th

Crew 210 Operations Report 15-APR-2019

SOL: 3

Name of person filing report: Shannon Rupert

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run): 11 hours last night

Solar— SOC 89 % as of COMMS. Will not turn on generator until needed.

Diesel Reading – 38 %

Propane Reading – 45 %

Ethanol Free Gasoline – 10 gallons.

Water (auxillary tank) – 0 gallons

Water (static tank) – 300 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – unknown gallons.

Water (loft) – Static to Loft Pump used – yes

Water Meter: not read

Toilet tank emptied: no

Deimos rover used: Still non-operational


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no

Hours: 50.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: No

Hours: 86.5

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: yes, to move water tank

Hours: 56.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: No

Hours: 88.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) 350.1, 350.2, 350.3

Reason for use: Testing

Oil Added? No

ATV Fuel Used: 0 Gals.

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.1 was the only operational ATV when I arrived.
350.2 and 350.3 are now operational with new batteries. 350.2 is
still very dirty.

We are currently changing the spark plug on the Honda to see if we can
get it to turn over. The Honda’s battery is good. We purchased a new
battery for the 300 and hope to get it installed before dark today.
The goal is to have all 5 ATVs operational tomorrow.

HabCar used and why, where? Yes, to haul water.

CrewCar used and why, where? Yes, to Loa for supplies

General notes and comments: Tunnel repairs were completed. A
considerable amount of MOOP was picked up from around the tunnels.

We are currently transferring water with a very small external pump
that takes over an hour to transfer 125 gallons. We plan to replace
the pump that was broken last month with my old house pump this
evening. That should make transferring water easier although we
should order a replacement pump to have on hand.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Electric kennel is broken.
Toaster/convection oven was replaced on the shelf where it belongs.

Summary of GreenHab operations: David will send a full GreenHab report
tonight. He has programmed the fan to keep the plants cooler in the
afternoon. He will continue fine tuning it for the rest of the week.
He also purchased a new thermometer.

Summary of ScienceDome operations: First group of plants in the plant
press are being fanned for more rapid drying.

Summary of RAM operations: Clean! I spent part of the afternoon
organizing and cleaning the RAM. Happy to say it is now clean, except
that I will sweep it out once the wind dies down.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing tonight,
thank you.

Sol Summary – April 15th

Crew 210 Sol Summary Report 15-APR-2019

Sol: 3

Summary Title: Squashing plants and supply runs

Author’s name: Paul Sokoloff

Mission Status: Nominal

Sol Activity Summary: This morning, after the requisite coffee, Crew 210 pressed the plants collected during yesterdays excursion to Salt Creek and the Burpee Quarry. By squishing the plant specimens between cardboard and drying them out, the team is left with a flat plant that, of stored correctly, will last for hundreds of years. These specimens, accompanied by a label indicating the location and date of collection, will serve as the proof that these plants were found growing at a specific place and time, and will make up the backbone of our first planned publication from this program.

Our team devoted the rest of this sol to gathering supplies essential to hab and rover systems. Shannon made a water run to Hanksville, while the rest of 210 went to Bicknel for ATV batteries and supplies to build a quadrat for our ecological work. On our return to MDRS our team was able to repair one more ATV, three of these are now operational.

While completing both our scientific and technical work for the day we were also able to continue our social outreach efforts, including posting video of our plant pressing adventures to Twitter and Facebook, preparing for an “Ask a Martian Anything” on Reddit tomorrow night, and telling our “Martian” tales on social media.

All in all, another beautiful day in that special place that straddles the line between our unique planet and the one we’d next like to visit.

Look Ahead Plan: Now equipped with a square-meter quadrat for ecological sampling, we will plan on spending tomorrow (sol 4), doing transects and quadrats within the MDRS operational area. At these sites we will document the number of species, percent cover, and functional groups of the plant and lichen communities, and record abiotic variables such as soil moisture content, pH, and coarseness. New plant and lichen species found on EVA will also be collected.

A Bureau of Land Management representative is scheduled to visit MDRS tomorrow morning, so we will meet with them before starting our sampling EVA.

Anomalies in work: No.

Weather: Cloudy in the am, clearing by mid-day.

Crew Physical Status: Good.

EVA: no EVAs at MDRS, but Shannon made a water and supply run to Hanksville, and the remainder of the crew made a supply run to Bicknell.

Reports to be filed:

Sol Summary

Operations Report

GreenHab Report

Photo Report

Support Requested: n/a

Copyright © The Mars Society. All rights reserved. | Main Site